A box slides down a ramp from a truck bed at an angle of 37.0 degrees with the horizontal. How far (horizontally) is the bottom of the ramp from the truck bed if the vertical height of the ramp is 6.00 m?
1
Expert's answer
2012-06-01T09:07:09-0400
A box slides down a ramp from a truck bed at an angle of 37.0 degrees
with
the horizontal. How far (horizontally) is the bottom of the ramp
from the
truck bed if the vertical height of the ramp is 6.00 m?
l = h / tan(37.0) =
6.00 / 0.7536 = 7.96 m